# Detection and Geographic Localization of Natural Objects in the Wild: A Case Study on Palms

### v1

@@ 5.3 Calibration Analysis and Saliency Maps
To enhance confidence reliability in palm detection, we evaluated calibration performance before and after applying calibration techniques. Confidence scores should align with IoU, and LRP-based thresholding ($A$) is used to filter out low-confidence detections [kuzucu2025calibration]. As shown in Table 2, IR achieves the best LaECE0 for most models, while YOLOv10 benefits most from TS and DDQ from LR. For LaACE0, the optimal method varies – TS improves RTDETR and YOLOv10, whereas LR is most effective for DDQ and YOLO11. Notably, no single method is universally best, and in some cases, lowering LaECE0 increases LaACE0. Figure 7 further illustrates the effect of LRP-based thresholding and calibration on YOLOv10’s confidence distribution relative to IoU. Initially, low-confidence predictions ($<0.05$) dominate, skewing calibration metrics. LRP filtering refines the distribution by removing unreliable detections. Post-calibration, both LaECE0 and LaACE0 further decrease, indicating improved alignment between confidence scores and IoUs.

### v2

@@ 5.3 Calibration Analysis and Saliency Maps
To enhance confidence reliability in palm detection, we evaluated calibration performance before and after applying calibration techniques. Confidence scores should align with IoU, and LRP-based thresholding ($A$) is used to filter out low-confidence detections [22]. As shown in Table 2, IR achieves the best LaECE0 for most models, while YOLOv10 benefits most from TS and DDQ from LR. For LaACE0, the optimal method varies – TS improves RTDETR and YOLOv10, whereas LR is most effective for DDQ and YOLO11. Notably, no single method is universally best, and in some cases, lowering LaECE0 increases LaACE0. Figure 7 further illustrates the effect of LRP-based thresholding and calibration on YOLOv10’s confidence distribution relative to IoU. Initially, low-confidence predictions (<0.05<0.05) dominate, skewing calibration metrics. LRP filtering refines the distribution by removing unreliable detections. Post-calibration, both LaECE0 and LaACE0 further decrease, indicating improved alignment between confidence scores and IoUs.